Using the allowance method of estimating bad debt, with the percent of sales, determine the amount to be added to “Allowance for Doubtful Account”, with the following information: Credit balance of $5,500 in Allowance for Doubtful Accounts just prior to adjustment. Uncollectible receivables are estimated at 2% of credit sales, which totaled $7,000,000 for the year. (10 points) a. Calculate the amount of bad debt? b. What will the balance be in the allowance account? c. Please journalize the year end adjusting entry
Answer #1

a) Bad debt expense = 7000000*2% = 140000

b) Balance in Allowance account = 140000+5500 = $145500

c) Adjusting entry

Date account and explanation Debit Credit
bad debt expense 140000
Allowance for doubtful accounts 140000